gcc8617:

whats is 34x10+{3x4)+7

Justyn:

352x + 7

XioGonz:

Alright well we must use the alternative approach

boredfr:

@justyn wrote:
352x + 7
I think "x" is the multiplication sign not variable

Justyn:

Oh is it?

Justyn:

LOL

XioGonz:

Which in this case Is Order of Operations

simsharrison:

use pemdas 0-0

XioGonz:

We begin by calculating the expression inside the parentheses. Perform any multiplication or division, from left to right SO

Justyn:

359

XioGonz:

\[34*10+(3*4)+7\]

XioGonz:

First comes Parenthesis SO 4*3

Justyn:

Parentheses (3*4), multiply and divide left to right 34*10 = 340, add and subtract left to right 340 + 12 + 7 = 359

XioGonz:

So we have \[34*10+(12)+7\]

XioGonz:

Now Perform any multiplication or division, from left to right:

boredfr:

First do the parenthesis (3x4)=12 34x10+12+7 next do multiplication 34x10= 340 so 340+12+7 340+12= 352 352+7= 359 Your answer is 359

Justyn:

You're slow Chad @xio

XioGonz:

\[34*10+12+7\] \[340+12+7\]

XioGonz:

Calculate any addition or subtraction, from left to right. \[340+\] \[352+7\] \[359\] So 359 is your final answer

XioGonz:

@boredfr wrote:
First do the parenthesis (3x4)=12 34x10+12+7 next do multiplication 34x10= 340 so 340+12+7 340+12= 352 352+7= 359 Your answer is 359
I've been warned for doing this before.

boredfr:

@xiogonz wrote:
@boredfr wrote:
First do the parenthesis (3x4)=12 34x10+12+7 next do multiplication 34x10= 340 so 340+12+7 340+12= 352 352+7= 359 Your answer is 359
I've been warned for doing this before.
Imagine-

XioGonz:

@boredfr wrote:
@xiogonz wrote:
@boredfr wrote:
First do the parenthesis (3x4)=12 34x10+12+7 next do multiplication 34x10= 340 so 340+12+7 340+12= 352 352+7= 359 Your answer is 359
I've been warned for doing this before.
Imagine-
frl though By Angle And Tranq Just ask lol

boredfr:

@xiogonz wrote:
@boredfr wrote:
@xiogonz wrote:
@boredfr wrote:
First do the parenthesis (3x4)=12 34x10+12+7 next do multiplication 34x10= 340 so 340+12+7 340+12= 352 352+7= 359 Your answer is 359
I've been warned for doing this before.
Imagine-
frl though By Angle And Tranq Just ask lol
Damn, well this is a simple equation so I don't plan on being like "What is 3 times 4 when it is basic math ._.

XioGonz:

Welp e.e